"\intop" 1/(3-2cos"\theta" +sin"\theta" ) using contour integration limit 0 to 2pi


1
Expert's answer
2022-01-31T16:43:52-0500

"1/(3-2cos\theta +sin\theta )=-\frac{2(1-tan^2(\theta /2))}{tan^2(\theta /2)+1}+\frac{2tan(\theta /2)}{tan^2(\theta /2)+1}+3"


"u=tan(\theta /2)\implies dx=2du/(u^2+1)"


"∫ 1/(3-2cos\theta +sin\theta)d\theta =2\int \frac{1}{2(u^2+u)+3u^2+1}du=2\int \frac{1}{(u\sqrt 5+1/\sqrt 5)^2+4/5}du="


"v=(5u+1)/2\implies du 2dv/5"


"=2\int \frac{2}{5(4v^2/5+4/5)}dv=\int \frac{dv}{1+v^2}=arctanv=arctan \frac{5u+1}{2}=arctan\frac{5tan(\theta/2)+1}{2}"


"∫^{2\pi}_0 1/(3-2cos\theta +sin\theta)d\theta=arctan\frac{5tan(\theta/2)+1}{2}|^{2\pi}_0="


"=arctan 1\/2-arctan 1\/2=0"